# Higher memory effects and the post-Newtonian calculation of their gravitational-wave signals

### v1

@@ 1 Introduction
While the three memory effects (displacement, CM and spin) have close connections with the symmetries and conserved quantities of the BMS framework and its extensions and generalizations, there are a large class of lasting gravitational-wave phenomena, called “persistent observables” by [20], which did not need to be computed in the context of asymptotically flat spacetimes and did not have obvious relationships to symmetries and conserved quantities. One of the observables of [20], the curve deviation, is a generalization of the (in-principle) procedure used to measure the memory effect. Specifically, it allows the test masses that are used to measure the relative separation to be in relative acceleration rather than freely falling, and they are no longer assumed to be initially comoving. The relative separation will change because of both the initial relative velocity and the relative acceleration, but once these kinematical effects are removed, there is a residual change in the final relative separation that depends on the spacetime curvature and its time integrals, as well as the initial separation, velocity, and acceleration, as well as all the initial time derivatives of the acceleration. This residual change was what was described as the curve deviation observable.
@@
These new moments of the news have flux and charge contributions, and we will focus on the flux contributions in this paper. For higher moments of the news, the flux terms involve an increasing number of integrals of fluxes that have appeared in lower moments of the news; throughout this paper, we will focus on the “new” parts of the fluxes, which do not appear in the expressions for lower moments of the news. These fluxes can be written in terms of products of the shear and news (and their angular derivatives), and we derive general multipolar expressions for these new fluxes in terms of a set of radiative multipole moments of the shear. Starting at the second moment of the news, we also compute the multipolar expansion of “pseudo-fluxes” (quantities that depend on the shear or other so-called “non-radiative” Bondi metric functions). We compute the fluxes and pseudo-fluxes necessary to determine the second moment of the news; however, our results could be extended to compute higher moments of the news.
@@
We also use the multipolar expansions to compute the leading PN expressions for the moments of the news (and their corresponding gravitational-wave signals) that are generated by the leading gravitational-wave strain produced during the inspiral of nonspinning compact objects on quasicircular orbits. We recover the leading-order non-oscillatory effects for the nonlinear displacement and spin memory effects given in Refs. [33] and [14], respectively, which are computed from the leading PN oscillatory waveforms without memory. We also compute “oscillatory memory” terms, which are nonlinear contributions to the oscillatory waveform modes of the strain that arise from the same flux and pseudo-flux terms used to compute the more familiar non-oscillatory memory effects. Our PN calculation of the oscillatory CM memory corrects the result in [15]; however, we also complete calculations of the non-oscillatory and oscillatory displacement, spin and CM memories that we believe were not computed previously. For the second moment of the news and the corresponding strain, the oscillatory memory terms prove to be larger than the non-oscillatory terms. Since the gravitational-wave signal (the strain as a function of time) is the observable that could be measured by gravitational-wave detectors, the oscillatory memory could be the easier to detect. However, because we compute the PN expressions during the inspiral only (we would need to use numerical-relativity waveforms to cover the merger and ringdown phases; see [34]), we do not give quantitative estimates of the detection prospects here, as they would be underestimates due to the truncation of the signal.
@@ 2 BMS formalism
The Bondi-Metzner-Sachs formalism (see, e.g., the review [35]) is frequently used to study asymptotically flat spacetimes, because the coordinates are well adapted to outgoing radiation. We denote our Bondi-Sachs coordinates by $x^{\alpha}=\{u,r,\theta^{A}\}$. The variable $u$ is the retarded time, and the $u=\mathrm{constant}$ surfaces are null (i.e., the dual vector $k_{\alpha}=-(\mathrm{d}u)_{\alpha}$ is null and normal to these hypersurfaces). The angular coordinates are represented here by $\theta^{A}$, with $A$, $B$ being coordinate indices on the two-sphere. These angular coordinates are also constant along the null rays to which $k^{\alpha}$ is tangent: $k^{\alpha}\partial_{\alpha}\theta^{A}=0$. Finally, $r$ is an areal radius, in the sense that the area of surfaces of constant $u$ and $r$ is proportional to $r^{2}$ and independent of $u$. The above constraints define Bondi gauge, and they imply
@@ 2.1 Metric, Einstein equations, boundary conditions, and expansion of the metric
The contracted Bianchi identity (specifically the $\beta=u$ and $\beta=A$ components of $\nabla_{\alpha}{R^{\alpha}}_{\beta}=0$ in vacuum) gives rise to the constraints $\partial_{r}(r^{2}e^{\frac{2\beta}{r}}R^{r}{}_{u})=0$ and $\partial_{r}(r^{2}e^{\frac{2\beta}{r}}R^{r}{}_{A})=0$, which imply that the $R^{r}{}_{u}$ and $R^{r}{}_{A}$ components of the vacuum Einstein equations are satisfied everywhere, if satisfied on any finite worldtube or in the limit $r\rightarrow\infty$. Thus $R^{r}{}_{u}=0$ and $R^{r}{}_{A}=0$ are known as the supplementary equations, as compared with the hypersurface and evolution equations, which together are referred to as the main equations. When the main equations are satisfied, the $\beta=r$ component of the Bianchi identity implies $(\mathcal{H}^{-1})^{AB}R_{AB}=0$, which shows that the trace part of $R_{AB}=0$ is not an independent equation. The main and supplementary equations then determine all of the metric functions in the Bondi gauge.
@@
Here, $m$ (the mass aspect) and $V^{A}$ are functions of integration of the hypersurface equations for $V$ and $\mathcal{U}^{A}$, respectively, and thus are functions of $u$ and $\theta^{A}$. The functions $\tilde{\beta}$, $\mathcal{M}$, and $\Upsilon^{A}$ are functions of all four coordinates, and they admit an expansion in $1/r$ starting at $O(1)$; the coefficients in this expansion can be written in terms of the expansion of $\mathcal{H}_{AB}$, its derivatives, and the other functions of integration.
@@ 2.2 Evolution equations for the expanded Bondi metric functions
One reason for this definition is that $\hat{N}^{A}$ simplifies the integrand for the “conserved quantity” conjugate to the Lorentz symmetries of the BMS algebra at null infinity, as defined by Wald and Zoupas [9] (see also [12]).222 Other definitions of these conserved quantities have been discussed in the literature [36], and there are various arguments for and against them [37, 38, 39]. This definition will also be useful for comparing our PN calculations that we will perform in Sec. 6 with prior PN calculations of the spin and CM memory effects.
@@
(*$(^{*}T)_{A_{1}\cdots A_{s}}\equiv\epsilon^{B}{}_{A_{1}}T_{BA_{2}\cdots A_{s}}.$ (2.12)
@@ 3.2 Computing the moments of the news and the corresponding shear
However, there are often scenarios (most frequently for analytical approximation methods) in which the equations for the moments of the news can be used to calculate a contribution to a moment from the changes in charges or fluxes that involve approximate gravitational waveforms that do not contain this moment (at a given order in a particular approximation method, for example). Some of these contexts include in post-Newtonian theory (see, e.g., [33, 7, 14, 36]), linear perturbations of black holes [41], and even some numerical-relativity waveforms computed using extrapolation methods at finite radii (see, e.g., [42, 14, 43, 44, 45, 34]). [^6] Numerical relativity calculations using Cauchy-characteristic evolution or matching do not have this property. In these cases, the moments of the news can be used as a check of the numerical accuracy of the solutions.
@@ 4.1 Zeroth and first fluxes
The zeroth flux can be computed by taking the $u$ derivative of the expression (4.2) to obtain the multipolar expansion of the news tensor. This multipolar expansion can then be substituted into Eq. (2.14a). After re-writing the electric- and magnetic-parity tensor harmonics in terms of spin-weighted harmonics using Eqs. (D), the scalar multipolar moment $(\mathcal{F}_{0})_{lm}$ of the flux can be written in terms of the integral of a product of three spin-weighted spherical harmonics. These integrals can be written as proportional to the product of two Clebsch-Gordan coefficients with several selection rules determining when these integrals are nonzero. These expressions arise because the product of two spin-weighted spherical harmonics can be written as a sum of spin-weighted spherical harmonics multiplied by products of Clebsch-Gordan coefficients. The relevant expression is given in Eq. (D.17). The result has been derived before in several different notations (e.g., [33, 14, 45]); our result is most similar to the expression for $\Delta\Phi_{lm}$ in [14], though we use the slightly more compact notation that we introduce below.
@@
With the notation set, we can describe more briefly how we obtain the multipolar expressions for the first fluxes $\mathcal{F}_{1}$ and $\mathcal{F}_{1}^{*}$. Starting from Eqs. (2.14b) and (2.14c), we then substitute in the multipolar expansion of $C_{AB}$ (and its $u$ derivative) in (4.2) and make use of the “raising” and “lowering” expressions for STF tensor harmonics in Eqs. (D.11) and (D), respectively. The result can be written as
@@ 5.2 Post-Newtonian expressions for the shear associated with moments of the news
In the remainder of this section, we present the results for the leading PN expressions for the the shear associated with the moments of the news in the quadrupole and octupole modes. We will use several of these results in Sec. 6 when we specialize to compact-binary sources (though we will compute additional modes there). We then describe our procedure to compare the PN expressions for the multipole moments of the strain (as given in Sec. 2.4.5 of [30]) with the expressions that we derive from taking derivatives of moments of the news.
@@ 5.2.1 Expressions for the quadrupole and octupole waveforms generated by fluxes and pseudo-fluxes
Because there are angular operators acting on the shear (on the left-hand side) and fluxes and pseudo-fluxes (on the right-hand side), it is convenient to expand the left-hand side in spherical harmonics (in addition to the right-hand side, which we already expanded in Sec. 4). For each spherical harmonic mode, inverting the angular operators to determine the moments of the shear then becomes an algebraic operation. Specifically, if we use $U^{\mathcal{F}}_{lm}$ and $V^{\mathcal{F}}_{lm}$ to denote the electric and magnetic parts of the shear that arise from the flux and pseudo-flux terms for a given spherical harmonic mode, then we can write the contribution from the zeroth, first and second moments relatively concisely. To do so, we apply Eq. (D) twice to the expansion of $C_{AB}$ in terms of tensor harmonics and use the facts that $\mathcal{D}_{1}=\mathscr{D}^{2}$, $\mathcal{D}_{2}=\mathscr{D}^{2}+2$ and $\mathscr{D}^{2}Y_{lm}=-l(l+1)Y_{lm}$. It then follows that the electric part of the shear, $U^{\mathcal{F}}_{lm}$, is given by
@@ 5.2.2 Comparison with PN waveform multipoles
With the PN quadrupole $\mathcal{U}_{ij}$ and octupole $\mathcal{U}_{ijk}$ written in the form in Eqs. (A) and (A), we can now more easily compare the different flux and pseudo-flux contributions in $U^{\mathcal{F}}_{2m}$ and $U^{\mathcal{F}}_{3m}$, which we can do separately for each flux. What now remains is to convert between our expressions involving scalar multipoles and the STF tensor multipoles used in PN theory. This will require evaluating the combination of Clebsch-Gordan coefficients that appears in the scalar-multipole expressions (which we do using Mathematica) and some of the tensors $\bar{\mathcal{Y}}_{L}^{lm}$, which are given in Eq. (2.12) of [47]. To make the comparison for the quadrupole, we take each unique product of multipoles on the right-hand sides of Eq. (5.2.1) and replace the scalar radiative mass and current multipoles with Eq. (5.10). We then multiply by $a_{l}Y_{lm}$ and sum over $m$, so that we obtain the contribution of this term to $\mathcal{U}^{\mathcal{F}}_{ij}n^{i}n^{j}$. We then compare this with the equivalent term (in terms of multipole order and number of time derivatives acting on the multipoles) in Eq. (A) by contracting the expression there with $n^{i}n^{j}$, so as to verify that the two terms are equal. By repeating this process for all the terms, we find that the expression for the quadrupole computed from Eq. (5.2.1) [using the moments in Eq. (5.2.1)] are equivalent to that in Eq. (A), up to the error terms involving $\mathrm{d}^{2}/\mathrm{d}u^{2}$ of products of canonical multipoles. We repeat the same process for the octupole.
@@
Substituting the expressions for the fluxes from Eq. (5.2.2) into Eq. (5.2.2) and comparing with Eq. (A), shows, in detail, how the PN expansion of the octupole moment can be understood in terms of the contributions from the flux and pseudo-flux pieces of the moments of the news. The PN memory integral comes from the zeroth flux $\mathcal{F}_{0}$, the instantaneous terms with no retarded-time derivatives come from the first flux $\mathcal{F}_{1}$ and both radiative and non-radiative the pseudo-flux pieces of the second moment contribute to the first total derivative term. The higher derivative terms could, in principle, be obtained from calculating flux and pseudo-flux contributions from the third moment of the news, although this still neglects the charge contribution to the third moment.
@@
This implies the the full flux contributions to the quadrupole can be written in STF form as
@@
Substituting the STF multipole expressions for the fluxes and pseudo-fluxes in Eq. (5.2.2) into Eq. (5.19) reproduces the waveform in Eq. (A) up to the second-derivative terms. As with the octupole, the PN memory integral arises from the zeroth flux, the instantaneous terms without derivatives come from the first flux, and the second fluxes and pseudo-fluxes contribute to the term involving a single time derivative. For the quadrupole case, the higher derivative terms cannot be obtained from flux and pseudo-flux contributions to higher moments of the news, because the angular operators which appear on left-hand side of Eq. (3.25) remove the $l=2$ component of the shear for $n\geq 3$. This follows from Eq. (2.17): the derivative operator $\mathcal{D}_{n}$ acting on $Q_{n-1}$ in the evolution equation for $Q_{n}$ sets the $l=n-1$ mode to zero. This is why we do not include a residual term in Eq. (5.19). There also will be charge contributions from the $n=2$ charge in the PN expression that we have been neglecting.

### v2

@@ 1 Introduction
While the three memory effects (displacement, CM and spin) have close connections with the symmetries and conserved quantities of the BMS framework and its extensions and generalizations, there are a large class of lasting gravitational-wave phenomena, called “persistent observables” by [20], which did not need to be computed in the context of asymptotically flat spacetimes and did not have obvious relationships to symmetries and conserved quantities. One of the observables of [20], the curve deviation, is a generalization of the (in-principle) procedure used to measure the memory effect. Specifically, it allows the test masses that are used to measure the relative separation to be in relative acceleration rather than freely falling, and they are no longer assumed to be initially comoving. The relative separation will change because of both the initial relative velocity and the relative acceleration, but once these kinematical effects are removed, there is a residual change in the final relative separation that depends on the spacetime curvature and its time integrals, as well as the initial separation, velocity, acceleration, and all initial time derivatives of the acceleration. This residual change was what was described as the curve deviation observable.
@@
These moments of the news have flux and charge contributions, and we will focus on the flux contributions in this paper. For higher moments of the news, the flux terms involve an increasing number of integrals of fluxes that have appeared in lower moments of the news; throughout this paper, we will focus on the “new” parts of the fluxes, which do not appear in the expressions for lower moments of the news. These fluxes can be written in terms of products of the shear and news (and their angular derivatives), and we derive general multipolar expressions for these new fluxes in terms of a set of radiative multipole moments of the shear. Starting at the second moment of the news, we also compute the multipolar expansion of “pseudo-fluxes” (quantities that depend on the shear or other so-called “non-radiative” Bondi metric functions). We compute the fluxes and pseudo-fluxes necessary to determine the second moment of the news; however, our results could be extended to compute higher moments of the news.
@@
We also use the multipolar expansions to compute the leading PN expressions for the moments of the news (and their corresponding gravitational-wave signals) that are generated by the leading gravitational-wave strain produced during the inspiral of nonspinning compact objects on quasicircular orbits. We recover the leading-order non-oscillatory effects for the nonlinear displacement and spin memory effects given in Refs. [33] and [14], respectively, which are computed from the leading PN oscillatory waveforms without memory. We also compute “oscillatory memory” terms, which are nonlinear contributions to the oscillatory waveform modes of the strain that arise from the same flux and pseudo-flux terms used to compute the more familiar non-oscillatory memory effects. Our PN calculation of the oscillatory CM memory corrects the result in [15]; however, we also complete calculations of the non-oscillatory and oscillatory displacement, spin and CM memories that we believe were not computed previously. For the second moment of the news and the corresponding strain, the oscillatory memory terms prove to be larger than the non-oscillatory terms. Since the gravitational-wave signal (the strain as a function of time) is the observable that could be measured by gravitational-wave detectors, the oscillatory memory could be the easier to detect. However, because we compute the PN expressions during the inspiral only (we would need to use numerical relativity waveforms to cover the merger and ringdown phases; see [34]), we do not give quantitative estimates of the detection prospects here, as they would be underestimates due to the truncation of the signal.
@@ 2 BMS formalism
The BMS formalism (see, e.g., the review [35]) is frequently used to study asymptotically flat spacetimes, because the coordinates are well adapted to outgoing radiation. We denote our Bondi-Sachs coordinates by $x^{\alpha}=\{u,r,\theta^{A}\}$. The variable $u$ is the retarded time, and the $u=\mathrm{constant}$ surfaces are null (i.e., the dual vector $k_{\alpha}=-(\mathrm{d}u)_{\alpha}$ is null and normal to these hypersurfaces). The angular coordinates are represented here by $\theta^{A}$, with $A$, $B$ being coordinate indices on the two-sphere. These angular coordinates are also constant along the null rays to which $k^{\alpha}$ is tangent: $k^{\alpha}\partial_{\alpha}\theta^{A}=0$. Finally, $r$ is an areal radius, in the sense that the area of surfaces of constant $u$ and $r$ is proportional to $r^{2}$ and independent of $u$. The above constraints define Bondi gauge, and they imply
@@ 2.1 Metric, Einstein equations, boundary conditions, and expansion of the metric
The contracted Bianchi identity (specifically the $\beta=u$ and $\beta=A$ components of $\nabla_{\alpha}{R^{\alpha}}_{\beta}=0$ in vacuum) gives rise to the constraints $\partial_{r}(r^{2}e^{\frac{2\beta}{r}}R^{r}{}_{u})=0$ and $\partial_{r}(r^{2}e^{\frac{2\beta}{r}}R^{r}{}_{A})=0$, which imply that the $R^{r}{}_{u}$ and $R^{r}{}_{A}$ components of the vacuum Einstein equations are satisfied everywhere, if they are satisfied on any finite worldtube or in the limit $r\rightarrow\infty$. Thus $R^{r}{}_{u}=0$ and $R^{r}{}_{A}=0$ are known as the supplementary equations, as compared with the hypersurface and evolution equations, which together are referred to as the main equations. When the main equations are satisfied, the $\beta=r$ component of the Bianchi identity implies $(\mathcal{H}^{-1})^{AB}R_{AB}=0$, which shows that the trace part of $R_{AB}=0$ is not an independent equation. The main and supplementary equations then determine all of the metric functions in the Bondi gauge.
@@
Here, $m$ (the mass aspect) and $V^{A}$ are functions of integration of the hypersurface equations for $V$ and $\mathcal{U}^{A}$, respectively, and thus are functions of $u$ and $\theta^{A}$. The functions $\tilde{\beta}$, $\mathcal{M}$, and $\Upsilon^{A}$ depend on all four coordinates, and they admit an expansion in $1/r$ starting at $O(1)$; the coefficients in this expansion can be written in terms of the expansion of $\mathcal{H}_{AB}$, its derivatives, and the other functions of integration.
@@ 2.2 Evolution equations for the expanded Bondi metric functions
One reason for this definition is that $\hat{N}^{A}$ simplifies the integrand for the “conserved quantity” conjugate to the Lorentz symmetries of the BMS algebra at null infinity, as defined by Wald and Zoupas [9] (see also [12]).333 Other definitions of these conserved quantities have been discussed in the literature [36], and there are various arguments for and against them [37, 38, 39]. This definition will also be useful for comparing our PN calculations that we will perform in Sec. 6 with prior PN calculations of the spin and CM memory effects.
@@
$(^{*}T)_{A_{1}\cdots A_{s}}\equiv\epsilon^{B}{}_{A_{1}}T_{BA_{2}\cdots A_{s}}.$ (2.12)
@@ 3.2 Computing the moments of the news and the corresponding shear
However, there are often scenarios (most frequently for analytical approximation methods) in which the equations for the moments of the news can be used to calculate a contribution to a moment from the changes in charges or fluxes that involve approximate gravitational waveforms that do not contain this moment (at a given order in a particular approximation method, for example). Some of these contexts include in post-Newtonian theory (see, e.g., [33, 7, 14, 36]), linear perturbations of black holes [41], and even some numerical relativity waveforms computed using extrapolation methods at finite radii (see, e.g., [42, 14, 43, 44, 45, 34]). [^7] Numerical relativity calculations using Cauchy-characteristic evolution or matching do not have this property. In these cases, the moments of the news can be used as a check of the numerical accuracy of the solutions.
@@ 4.1 Zeroth and first fluxes
The zeroth flux can be computed by taking the $u$ derivative of the expression (4.2) to obtain the multipolar expansion of the news tensor. This multipolar expansion can then be substituted into Eq. (2.14a). After re-writing the electric- and magnetic-parity tensor harmonics in terms of spin-weighted harmonics using Eqs. (D.6), the scalar multipolar moment $(\mathcal{F}_{0})_{lm}$ of the flux can be written in terms of the integral of a product of three spin-weighted spherical harmonics. These integrals can be written as proportional to the product of two Clebsch-Gordan coefficients with several selection rules determining when these integrals are nonzero. These expressions arise because the product of two spin-weighted spherical harmonics can be written as a sum of spin-weighted spherical harmonics multiplied by products of Clebsch-Gordan coefficients. The relevant expression is given in Eq. (D.17). The result has been derived before in several different notations (e.g., [33, 14, 45]); our result is most similar to the expression for $\Delta\Phi_{lm}$ in [14], though we use the slightly more compact notation that we introduce below.
@@
With the notation set, we can describe more briefly how we obtain the multipolar expressions for the first fluxes $\mathcal{F}_{1}$ and $\mathcal{F}_{1}^{*}$. Starting from Eqs. (2.14b) and (2.14c), we then substitute in the multipolar expansion of $C_{AB}$ (and its $u$ derivative) in (4.2) and make use of the “raising” and “lowering” expressions for STF tensor harmonics in Eqs. (D.11) and (D.13), respectively. The result can be written as
@@ 5.1 Relation between post-Newtonian and Bondi-Sachs waveforms
One can also compute the polarizations $h_{+}$ and $h_{\times}$ via the relationship
$h\equiv h_{+}-ih_{\times}={\sum}_{l,m}h_{lm}(_{-2}Y_{lm}),$ (5.5)
where the moments $h_{lm}$ are related to $U_{lm}$ and $V_{lm}$ by
$h_{lm}=\frac{1}{r\sqrt{2}}(U_{lm}-iV_{lm}),$ (5.6)
and where ${}_{-2}Y_{lm}$ is a spin-weighted spherical harmonic. The conventions that we use for these harmonics are defined in D.
@@ 5.2 Post-Newtonian expressions for the shear associated with moments of the news
In the remainder of this section, we present the results for the leading PN expressions for the shear associated with the moments of the news in the quadrupole and octupole modes. We will use several of these results in Sec. 6 when we specialize to compact-binary sources (though we will compute additional modes there). We then describe our procedure to compare the PN expressions for the multipole moments of the strain (as given in Sec. 2.4.5 of [30]) with the expressions that we derive from taking derivatives of moments of the news.
@@ 5.2.1 Expressions for the quadrupole and octupole waveforms generated by fluxes and pseudo-fluxes
Because there are angular operators acting on the shear (on the left-hand side) and fluxes and pseudo-fluxes (on the right-hand side), it is convenient to expand the left-hand side in spherical harmonics (in addition to the right-hand side, which we already expanded in Sec. 4). For each spherical harmonic mode, inverting the angular operators to determine the moments of the shear then becomes an algebraic operation. Specifically, if we use $U^{\mathcal{F}}_{lm}$ and $V^{\mathcal{F}}_{lm}$ to denote the electric and magnetic parts of the shear that arise from the flux and pseudo-flux terms for a given spherical harmonic mode, then we can write the contribution from the zeroth, first and second moments relatively concisely. To do so, we apply Eq. (D.13) twice to the expansion of $C_{AB}$ in terms of tensor harmonics and use the facts that $\mathcal{D}_{1}=\mathscr{D}^{2}$, $\mathcal{D}_{2}=\mathscr{D}^{2}+2$ and $\mathscr{D}^{2}Y_{lm}=-l(l+1)Y_{lm}$. It then follows that the electric part of the shear, $U^{\mathcal{F}}_{lm}$, is given by
@@ 5.2.2 Comparison with PN waveform multipoles
With the PN quadrupole $\mathcal{U}_{ij}$ and octupole $\mathcal{U}_{ijk}$ written in the form in Eqs. (A) and (A), we can now more easily compare the different flux and pseudo-flux contributions in $U^{\mathcal{F}}_{2m}$ and $U^{\mathcal{F}}_{3m}$, which we can do separately for each flux. What now remains is to convert between our expressions involving scalar multipoles and the STF tensor multipoles used in PN theory. This will require evaluating the combination of Clebsch-Gordan coefficients that appears in the scalar-multipole expressions (which we do using Mathematica) and some of the tensors $\bar{\mathcal{Y}}_{L}^{lm}$, which are given in Eq. (2.12) of [47]. To make the comparison for the quadrupole, we take each unique product of multipoles on the right-hand sides of Eq. (5.16) and replace the scalar radiative mass and current multipoles with Eq. (5.12). We then multiply by $a_{l}Y_{lm}$ and sum over $m$, so that we obtain the contribution of this term to $\mathcal{U}^{\mathcal{F}}_{ij}n^{i}n^{j}$. We then compare this with the equivalent term (in terms of multipole order and number of time derivatives acting on the multipoles) in Eq. (A) by contracting the expression there with $n^{i}n^{j}$, so as to verify that the two terms are equal. By repeating this process for all the terms, we find that the expression for the quadrupole computed from Eq. (5.2.1) [using the moments in Eq. (5.16)] are equivalent to that in Eq. (A), up to the error terms involving $\mathrm{d}^{2}/\mathrm{d}u^{2}$ of products of canonical multipoles. We repeat the same process for the octupole.
@@
Substituting the expressions for the fluxes from Eq. (5.17) into Eq. (5.2.2) and comparing with Eq. (A), shows, in detail, how the PN expansion of the octupole moment can be understood in terms of the contributions from the flux and pseudo-flux pieces of the moments of the news. The PN memory integral comes from the zeroth flux $\mathcal{F}_{0}$, the instantaneous terms with no retarded-time derivatives come from the first flux $\mathcal{F}_{1}$ and both radiative and non-radiative the pseudo-flux pieces of the second moment contribute to the first total derivative term. The higher derivative terms could, in principle, be obtained from calculating flux and pseudo-flux contributions from the third moment of the news, although this still neglects the charge contribution to the third moment.
@@
This implies the full flux contributions to the quadrupole can be written in STF form as
@@
Substituting the STF multipole expressions for the fluxes and pseudo-fluxes in Eq. (5.20) into Eq. (5.21) reproduces the waveform in Eq. (A) up to the second-derivative terms. As with the octupole, the PN memory integral arises from the zeroth flux, the instantaneous terms without derivatives come from the first flux, and the second fluxes and pseudo-fluxes contribute to the term involving a single time derivative. For the quadrupole case, the higher derivative terms cannot be obtained from flux and pseudo-flux contributions to higher moments of the news, because the angular operators which appear on left-hand side of Eq. (3.25) remove the $l=2$ component of the shear for $n\geq 3$. This follows from Eq. (2.17): the derivative operator $\mathcal{D}_{n}$ acting on $Q_{n-1}$ in the evolution equation for $Q_{n}$ sets the $l=n-1$ mode to zero. This is why we do not include a residual term in Eq. (5.21). There also will be charge contributions from the $n=2$ charge in the PN expression that we have been neglecting.